# HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometry

### v1

@@ 4 HyperAIRI
4.1 Hyperspectral PnP-FB and HyperAIRI denoiser
By definition, the proximal operator associated with a regularisation function (7) can be interpreted as solving a denoising problem. PnP algorithms build on this insight by replacing the proximal operator with a pretrained DNN, typically trained for image denoising (Venkatakrishnan et al., 2013). In the case of additive Gaussian noise, the denoising problem consists of estimating the ground truth $\overline{\mathbf{U}}$ from the noisy measurement $\mathbf{Z}$, modelled as
@@ 4.1 Hyperspectral PnP-FB and HyperAIRI denoiser
where $\mathbf{E}\sim\mathcal{N}(0,\mathbf{I})$ is a realisation of a white Gaussian noise cube with standard deviation and $\sigma$. By construction, the trained denoisers are blind to the measurement process, and can therefore adapt easily to RI imaging for different telescopes and under varying observational conditions. AIRI (Terris et al., 2023, 2025) is a successful instantiation of PnP algorithms for monochromatic RI imaging. By replacing the handcrafted proximity operators in FB with learned firmly non-expansive denoisers [^4] An operator $\mathrm{D}(\cdot)$ is firmly non-expansive if it satisfies the condition: $\|\mathrm{D}(\bm{a})-\mathrm{D}(\bm{b})\|^{2}\leq\langle\mathrm{D}(\bm{a})-\mathrm{D}(\bm{b}),\bm{a}-\bm{b}\rangle$. Intuitively, this means $\mathrm{D}(\cdot)$ moves inputs closer together while generally preserving their relative directions., AIRI provides state-of-the-art imaging precision, preserving the robustness of optimisation algorithms and the flexibility of PnP algorithms. Building on the monochromatic AIRI design, HyperAIRI adopts the PnP-variant of (6), which reads
$\mathbf{X}^{i+1}=\mathrm{D}\left(\mathbf{X}^{i}-\gamma\nabla f\left(\mathbf{X}^{i}\right)\right),$ (17)
where $\mathrm{D}(\cdot)$ represents denoising via a trained DNN denoiser.
Furthermore, we adapt AIRI’s network architecture to accommodate multi-channel input and exploit spectral correlations. In hyperspectral RI imaging, the number of spectral channels can vary widely. Applying the DNN along the spectral axis via an overlapping window provides a flexible approach, decoupling the number of DNN input channels from the size of the target image cube during image reconstruction. This design also supports channel-wise parallelisation when denoising, albeit each denoiser only leverages correlations within its immediate neighbours. To mitigate the potential sub-optimality of this local joint denoising, the power-law spectral model is introduced to propagate global spectral information.
The architecture of a HyperAIRI denoiser takes a variation of the DnCNN for AIRI containing 20 convolutional layers. To denoise a channel $l$ of a noisy image cube $\mathbf{Z}$, HyperAIRI denoiser takes as input (i) the sub-cube composed of the target channel $\bm{z}_{l}$ and its two adjacent channels, (ii) the frequency ratios between the neighbouring and target channels, and (iii) the estimated spectral index map. It then outputs the denoised image $\widetilde{\bm{z}}_{l}$. HyperAIRI denoiser injects the power-law spectral model underpinning the hyperspectral image cube via a spectral correction layer shown in Fig. 2(b). This layer takes the same set of inputs of HyperAIRI denoisers, providing as output the estimate of the target channel:
@@
where $\breve{\nu}_{l-1,l}=\nu_{l-1}/\nu_{l}$ and $\breve{\nu}_{l+1,l}=\nu_{l+1}/\nu_{l}$. In the spectral correction layer, only the spectral index map is involved, while the spectral curvature map is delegated implicitly to the other layers. This separation supports the numerical stability of the entire algorithm, especially in cases where $\widetilde{\bm{\alpha}}$ is estimated dynamically during reconstruction. The remaining layers in the DNN are designed to predict the residual difference between the initial estimate and the ground-truth image. The final estimate of the HyperAIRI denoiser is the summation of the predicted difference and the output of the spectral correlation layer with positivity constraint from ReLU layer. The full architecture is shown in Fig. 2(a).
As HyperAIRI denoiser targets single-channel output with similar numbers of learnable parameters, the training and inference costs are comparable to those of AIRI, maintaining its computational efficiency. Moreover, this sliding-window denoising strategy provides the flexibility to process image cubes of varying sizes. Thus, the proposed HyperAIRI denoiser structure meets our requirements for efficiency and flexibility.
4.2 Training HyperAIRI denoisers
To address the limited availability of training data for hyperspectral RI imaging, we build a training dataset based on the two monochromatic datasets proposed by Terris et al. (2025), namely the Optical Astronomical Image Dataset (OAID) and the Magnetic Resonance Image Dataset (MRID). Both datasets contain 5,000 images of size $512\times 512$ pixels, and are used jointly during training. For each monochromatic image, a corresponding hyperspectral image cube is generated according to the power-law model in (1). Following Thouvenin et al. (2023a), the curvature map $\bm{\beta}$ is modelled as a random Gaussian field, while the spectral index map $\bm{\alpha}$ is derived from the logarithm of a Gaussian-blurred version of $\overline{\bm{u}}$ combined with a random Gaussian field ensuring spatial correlation. The elements in $\bm{\alpha}$ are in the range [5,0][-5,0] and those in $\bm{\beta}$ are in the range [0.5,0.5][-0.5,0.5]. The simulated image cube $\overline{\mathbf{U}}\in\mathbb{R}^{N\times 3}$ is expressed as:
@@ 4.2 Training HyperAIRI denoisers
HyperAIRI denoisers are trained in a supervised manner using pairs of the simulated groundtruth image cubes $\overline{\mathbf{U}}$ and their noisy versions $\mathbf{Z}$ obtained following (16). As indicated by Terris et al. (2023, 2025), ensuring the non-expansiveness of the denoisers is crucial for the convergence of the final PnP algorithm in reconstructing HDR RI images. Therefore, the loss function includes a standard $\ell_{1}$ loss and a Lipschitz regularisation term. Denoting the denoiser as $\mathrm{D}$ with learnable parameters as $\bm{\theta}$, the training loss is
$\mathcal{L}(\bm{\theta},$ $\mathbf{Z},\mathbf{\overline{U}},\bm{\alpha},\breve{\bm{\nu}}_{2})=$ (20)
@@
$\|\operatorname{D}_{\bm{\theta}}(\mathbf{Z},\bm{\alpha},\breve{\bm{\nu}}_{2})-(\overline{\mathbf{U}})_{2}\|_{1}+\kappa\,\mathrm{max}\{\|\bm{\nabla}\operatorname{Q}_{\bm{\theta}}(\overline{\mathbf{U}})\|_{\rm{S}},1-\delta\},$
@@
where $(\cdot)_{2}$ selects the second channel of the image cube, $\bm{\alpha}$ and $\breve{\bm{\nu}}_{2}=(\breve{\nu}_{1,2},\breve{\nu}_{3,2})$ are the spectral index map and frequency ratios for $\mathbf{U}$. In the regularisation term, $\mathrm{Q}$ is defined as $2\mathrm{D}_{\bm{\theta}}(\cdot,\bm{\alpha},\breve{\bm{\nu}}_{2})-(\cdot)_{2}$, and $\bm{\nabla}\operatorname{Q}(\cdot)$ denotes the Jacobian of $\operatorname{Q}$. Here, $\kappa>0$ is the regularisation parameter, and $\delta>0$ determines the lower bound of the regularisation term.
@@
To avoid retraining for each specific problem, a shelf of HyperAIRI denoisers can be trained across various $\sigma$ values. In this work, we trained 6 DNNs at noise levels in the set $\{1.0\times 10^{-5},2.0\times 10^{-5},4.0\times 10^{-5},8.0\times 10^{-5},1.6\times 10^{-4},3.2\times 10^{-4}\}$ respectively, forming a shelf of HyperAIRI denoisers. Accompanied with proper scaling factors adaptively chosen during reconstructions, the denoisers will cover a dynamic range interval of $1.6\times 10^{3},1.0\times 10^{5}]$.
@@
Similarly to AIRI denoisers, HyperAIRI denoisers are trained using randomly cropped patches of size $46\times 46\times 3$ from the simulated image cubes to speed up the training process owing to their convolutional nature. The compact receptive fields allows application to any image size during image reconstruction.
@@
The training scripts are built with PyTorch in Python. Each HyperAIRI denoiser is trained for 5000 epochs with the ADAM optimiser (Kingma & Ba, 2014). The initial learning rate is set to $10^{-4}$, $10^{-5}$ and $10^{-6}$ for training noise levels $3.2\times 10^{-4}$, $4.0\times 10^{-5},8.0\times 10^{-5},1.6\times 10^{-4}]$ and $1.0\times 10^{-5},2.0\times 10^{-5}]$ respectively. The learning rate is halved every 900 epochs. The regularisation factor $\kappa$ in (20) is also fine-tuned for each DNN to make sure the final value of the regularisation term is as close to 1 as possible. In practice, $\kappa$ is typically close to the average value of the $\ell_{1}$ loss term for optimal results. As suggested by Terris et al. (2023), we also use the weights of the DNNs trained with $\kappa=0$ as initialisation to improve the training stability and the performance of the final DNNs.
4.3 Algorithmic structure
@@
Plugging HyperAIRI denoisers trained at desired noise levels into the iterative FB forms the HyperAIRI algorithm. The detailed structure of our proposed algorithm is shown in Fig. 2.(c).
4.3.1 Initialisation
@@
In the initialisation stage, each channel selects an appropriate denoiser from the pre-trained denoiser shelf based on its dynamic range. As suggested by Terris et al. (2023, 2025), the inverse of the target dynamic range at channel $l$ can be estimated as $\sigma_{\mathrm{heu},l}/\widetilde{\rho}_{l}$, where $\sigma_{\mathrm{heu},l}$ is given in (14), and $\widetilde{\rho}_{l}$ is the estimated maximum intensity of the image at channel $l$. As an initial estimation of $\widetilde{\rho}_{l}$, we consider the maximum intensity of the dirty image normalised by the peak value of the point spread function (PSF), i.e.
@@
$\widetilde{\rho}_{l}=\left.\max\left(\bm{x}_{l}^{\mathrm{dirty}}\right)\middle/\max\left(\mathrm{Re}\{\bm{\Phi}_{l}^{\dagger}\bm{\Phi}_{l}\bm{\delta}\}\right)\right$. (21)
@@
where $\bm{\delta}\in\mathbb{R}^{N}$ is a Dirac delta image with value 1 at its centre and 0 otherwise. For a given channel $l$, the denoiser with noise level $\sigma$ closest to but smaller than $\sigma_{\mathrm{heu},l}/\widetilde{\rho}_{l}$ is chosen. To adjust for the difference between the inverse dynamic range of the image and the noise level of the selected denoiser, a scaling factor is applied (Dabbech et al., 2022):
@@
with $\varrho_{l}=\sigma_{\mathrm{heu},l}/\sigma$. Finally, the spectral index map $\widetilde{\bm{\alpha}}$ can be initialised to $\bm{0}$, derived from the dirty image, or specified by the user.
4.3.2 Reconstruction loop
At each iteration $i$, the process for channel $l$ sends its gradient descent image $\hat{\bm{x}}^{i}_{l}$ to the adjacent channels (l1l-1 and $l+1$) and receives their corresponding images. These three images are combined to form the input cube $\hat{\mathbf{X}}^{i}_{l}=(\hat{\bm{x}}^{i}_{l-1},\hat{\bm{x}}^{i}_{l},\hat{\bm{x}}^{i}_{l+1})$, which is then processed by the denoiser for channel $l$. Specific to edge channels, the processes for the first and last channels form image cubes by communicating with their only neighbour, resulting in $\hat{\mathbf{X}}^{i}_{1}=(\hat{\bm{x}}^{i}_{1},\hat{\bm{x}}^{i}_{1},\hat{\bm{x}}^{i}_{2})$ and $\hat{\mathbf{X}}^{i}_{L}=(\hat{\bm{x}}^{i}_{L-1},\hat{\bm{x}}^{i}_{L},\hat{\bm{x}}^{i}_{L})$ respectively. Thereafter, the denoisers for each channel are applied simultaneously to $\hat{\mathbf{X}}^{i}_{l}$. Random $90^{\circ}$ rotations and flips, and the corresponding inverse transformations, are applied to the inputs and outputs of the DNNs, respectively, to exploit the transformation invariance of images (Terris et al., 2023). The result of the denoising step of all channels forms the image cube update $\mathbf{X}^{i+1}_{l}$.
4.3.3 Dynamic denoiser and spectral index map update
Following the standard FB update, the algorithm also updates the denoisers and the spectral index map dynamically. Initially, $\mathrm{D}_{l}$ is selected based on $\sigma_{\mathrm{heu},l}$ using the maximum intensity of the normalised dirty image as $\widetilde{\rho}_{l}$. However, this is a loose upper bound estimation of the real peak value $\rho_{l}$. To refine the estimate of $\sigma_{\mathrm{heu},l}$, $\widetilde{\rho}_{l}$ is updated as the maximum intensity of ${\bm{x}}^{i+1}_{l}$. $\sigma_{\mathrm{heu},l}$ is monitored at each iteration to determine whether the current denoiser $\operatorname{D}_{l}$ and the corresponding scaling factor $\varrho_{l}$ should be updated. Terris et al. (2025) has proven the effectiveness of this adaptive denoiser updating framework in monochromatic imaging. We adopt a similar strategy to update the estimate of the spectral index map $\widetilde{\bm{\alpha}}$ when needed, using the current estimate of the cube $\mathbf{X}^{i}$. Ignoring the spectral curvature term in (1), setting $\nu_{0}=1$ for simplicity and applying logarithm to both sides, the pixel $n$ in channel $l$ should follow
@@ 4.3.3 Dynamic denoiser and spectral index map update
Then, a linear relationship between $\log{\bm{x}_{l}}$ and $\log{\nu_{l}}$ is anticipated. The slope of the linear function can be evaluated pixel-wise using linear regression, which yields $\widetilde{\bm{\alpha}}$, the estimated spectral index map. In practice, a smoothing 2D Gaussian kernel is applied to the image cube before gathering them for linear regression to ensure numerical stability. The 2D Gaussian kernel is fitted to the PSF’s primary lobe at the highest-frequency channel, which represents the highest spatial resolution of the measurements. Futhermore, pixels with intensities below a preset threshold, taken as the minimal heuristic noise level of the entire image cube, are excluded from the estimation of $\widetilde{\bm{\alpha}}$.
4.3.4 Fully parallelised implementation
The entire HyperAIRI reconstruction is executed in a distributed, channel-wise fashion, together with an efficient large-scale RI measurement operator implementation (Dabbech et al., 2025). During initialisation, each process is assigned its channel’s back-projected dirty image, local measurement operator, selected denoiser and initial spectral-index map. At each FB iteration (see Fig. 2(d)), gradient-descent and denoising steps then run concurrently across channels. Each channel process exchanges only its immediate neighbours’ images to assemble the three-slice input cubes between these two steps, minimising communication overhead. As with the AIRI denoiser, the HyperAIRI denoiser has spatial faceting functionality. The denoisers can be applied on the overlapped facets extracted from $\hat{\mathbf{X}}^{i}_{l}$ in parallel. This hybrid spatial-spectral parallelisation scales nearly linearly with both the number of channels and available compute resources. As a result, our implementation delivers a highly efficient solver capable of handling large-scale hyperspectral datasets in practical runtimes.
5 Validation on simulated VLA observations

### v2

@@ IV HyperAIRI
IV.1 Hyperspectral PnP-FB and HyperAIRI denoiser
definition (7), the proximal operator associated with a regularization function can be interpreted as solving a denoising problem. PnP algorithms build on this insight by replacing the proximal operator with a pre-trained DNN, typically trained for image denoising (S. V. Venkatakrishnan et al., 2013). In the case of additive Gaussian noise, the denoising problem consists of estimating the ground truth $\overline{\mathbf{U}}$ from the noisy measurement $\mathbf{Z}$, modeled as
@@ IV.1 Hyperspectral PnP-FB and HyperAIRI denoiser
where $\mathbf{E}\sim\mathcal{N}(0,\mathbf{I})$ is a realization of a white Gaussian noise cube with standard deviation $\sigma$. By construction, the trained denoisers are blind to the measurement process, they can adapt to RI imaging for different telescopes and under varying observational conditions. AIRI has been a successful instantiation of PnP algorithms for monochromatic RI imaging. By replacing the handcrafted proximity operator in FB with learned firmly non-expansive denoiser [^4] An operator $\mathrm{D}(\cdot)$ is firmly non-expansive if it satisfies the condition: $\|\mathrm{D}(\bm{a})-\mathrm{D}(\bm{b})\|^{2}\leq\langle\mathrm{D}(\bm{a})-\mathrm{D}(\bm{b}),\bm{a}-\bm{b}\rangle$. Intuitively, this means $\mathrm{D}(\cdot)$ moves inputs closer together while generally preserving their relative directions., the robustness of optimization algorithms and the flexibility of PnP algorithms.
$\mathbf{X}^{i+1}=\mathrm{D}\left(\mathbf{X}^{i}-\gamma\nabla f\left(\mathbf{X}^{i}\right)\right).$ (17)
the HyperAIRI denoiser takes a DnCNN architecture with 20 convolutional layers. To denoise a channel $l$ of a noisy image cube $\mathbf{Z}$, the denoiser takes as input (i) the sub-cube composed of the target channel $\bm{z}_{l}$ and its two adjacent channels, (ii) the frequency ratios between the neighboring and target channels, and (iii) the estimated spectral index map. It then outputs the denoised image $\widetilde{\bm{z}}_{l}$. The HyperAIRI denoiser embeds the power-law spectral model underpinning the hyperspectral image cube via a spectral correction layer shown in Figure 2(b). This layer takes the same set of inputs of HyperAIRI denoisers, and provides as output the estimate of the target channel:
@@
where $\breve{\nu}_{l-1,l}=\nu_{l-1}/\nu_{l}$ and $\breve{\nu}_{l+1,l}=\nu_{l+1}/\nu_{l}$. In the spectral correction layer, only the spectral index map is involved, while the spectral curvature map is delegated implicitly to the other layers. This separation supports the numerical stability of the entire algorithm, especially in cases where $\widetilde{\bm{\alpha}}$ is estimated dynamically during reconstruction. The remaining layers in the DNN are designed to predict the residual difference between the initial estimate and the ground-truth image. The final estimate of the HyperAIRI denoiser is the summation of the predicted difference and the output of the spectral correlation layer with positivity constraint from ReLU layer. The full architecture is shown in Figure 2(a).
IV.2 Training HyperAIRI denoisers
To address the limited availability of training data for hyperspectral RI imaging, we build a training dataset based on the two monochromatic datasets proposed by M. Terris et al. (2025), namely the Optical Astronomical Image Dataset (OAID) and the Magnetic Resonance Image Dataset (MRID). Both datasets contain 5,000 images of size $512\times 512$ pixels, and are used jointly during training. For each monochromatic image, a corresponding hyperspectral image cube is generated according to the power-law model in (1). Following P.-A. Thouvenin et al. (2023a), the curvature map $\bm{\beta}$ is modeled as a random Gaussian field, while the spectral index map $\bm{\alpha}$ is derived from the logarithm of a Gaussian-blurred version of $\overline{\bm{u}}$ combined with a random Gaussian field ensuring spatial correlation. The elements in $\bm{\alpha}$ are in the range [5,0][-5,0] and those in $\bm{\beta}$ are in the range [0.5,0.5][-0.5,0.5]. The simulated image cube $\overline{\mathbf{U}}\in\mathbb{R}^{N\times 3}$ is expressed as:
@@ IV.2 Training HyperAIRI denoisers
HyperAIRI denoisers are trained in a supervised manner using pairs of the simulated ground-truth image cubes $\overline{\mathbf{U}}$ and their noisy versions $\mathbf{Z}$ obtained following (16). As indicated by M. Terris et al. (2023, 2025), ensuring the non-expansiveness of the denoisers is crucial for the convergence of the final PnP algorithm in reconstructing HDR RI images. Therefore, the loss function includes a standard $\ell_{1}$ loss and a Lipschitz regularization term. Denoting the denoiser as $\mathrm{D}$ with learnable parameters as $\bm{\theta}$, the training loss is
$\mathcal{L}(\bm{\theta},\mathbf{Z},\mathbf{\overline{U}},\bm{\alpha},\breve{\bm{\nu}}_{2})$ $=\|\operatorname{D}_{\bm{\theta}}(\mathbf{Z},\bm{\alpha},\breve{\bm{\nu}}_{2})-(\overline{\mathbf{U}})_{2}\|_{1}$ (20)
@@
$+$ $\kappa\,\mathrm{max}\{\|\bm{\nabla}\operatorname{Q}_{\bm{\theta}}(\overline{\mathbf{U}})\|_{\rm{S}},1-\delta\},$
@@
where $(\cdot)_{2}$ selects the second channel of the image cube, $\bm{\alpha}$ and $\breve{\bm{\nu}}_{2}=(\breve{\nu}_{1,2},\breve{\nu}_{3,2})$ are the spectral index map and frequency ratios for $\mathbf{U}$. In the regularization term, $\mathrm{Q}$ is defined as $2\mathrm{D}_{\bm{\theta}}(\cdot,\bm{\alpha},\breve{\bm{\nu}}_{2})-(\cdot)_{2}$, and $\bm{\nabla}\operatorname{Q}(\cdot)$ denotes the Jacobian of $\operatorname{Q}$. Here, $\kappa>0$ is the regularization parameter, and $\delta>0$ determines the lower bound of the regularization term.
@@
To avoid retraining for each specific problem, a shelf of HyperAIRI denoisers can be trained across various $\sigma$ values. In this work, we trained 6 DNNs at noise levels in the set
@@
Similarly to AIRI denoisers, HyperAIRI denoisers are trained using randomly cropped patches of size $46\times 46\times 3$ from the simulated image cubes to speed up the training process owing to their convolutional nature. The compact receptive field allows application to any image size during image reconstruction.
@@
The training scripts are built with PyTorch in Python. Each HyperAIRI denoiser is trained for 5000 epochs with the ADAM optimizer (D. P. Kingma & J. Ba, 2014). The initial learning rate is set to $10^{-4}$, $10^{-5}$ and $10^{-6}$ for training noise levels respectively. The learning rate is halved every 900 epochs. The regularization factor $\kappa$ in (20) is also fine-tuned for each DNN to make sure the final value of the regularization term is as close to 1 as possible. In practice, $\kappa$ is typically close to the average value of the $\ell_{1}$ loss term for optimal results. As suggested by M. Terris et al. (2023), we also use the weights of the DNNs trained with $\kappa=0$ as initialization to improve the training stability and the performance of the final DNNs.
IV.3 Algorithmic structure
@@
Plugging HyperAIRI denoisers trained at desired noise levels into the iterative FB forms the HyperAIRI algorithm. The detailed structure of our proposed algorithm is shown in Figure 2.(c).
IV.3.1 Initialization
@@
In the initialization stage, denoiser for each channel is selected from the pre-trained denoiser shelf on its dynamic range. As proposed by M. Terris et al. (2023), , where $\sigma_{\mathrm{heu},l}$ is the estimate of the image-domain noise level given in (14), and $\widetilde{\rho}_{l}$ is image at channel $l$. An initial estimate of $\widetilde{\rho}_{l}$ is obtained from the peak value of the dirty image normalized by the peak value of the point spread function (PSF), such that
@@
$\widetilde{\rho}_{l}=\left.\max\left(\bm{x}_{l}^{\mathrm{dirty}}\right)\middle/\max\left(\mathrm{Re}\{\bm{\Phi}_{l}^{\dagger}\bm{\Phi}_{l}\bm{\delta}\}\right)\right$., (21)
@@
where $\bm{\delta}\in\mathbb{R}^{N}$ is a Dirac delta image with value 1 at its center and 0 otherwise. For a given channel $l$, the denoiser with a noise level $\sigma$ closest to but smaller the target dynamic range estimate $\sigma_{\mathrm{heu},l}/\widetilde{\rho}_{l}$ is selected. To the noise level of the selected denoiser, a scaling factor is applied
@@
with $\varrho_{l}=\sigma_{\mathrm{heu},l}/\sigma$ (A. Dabbech et al., 2022). Finally, the spectral index map $\widetilde{\bm{\alpha}}$ can be initialized to $\bm{0}$, derived from the dirty image, or specified by the user.
IV.3.2
. At iteration $i$, the process assigned to each channel $l$ to the adjacent channels (l1l-1 and $l+1$) and receives their corresponding images. These three images are combined to form the input $\hat{\mathbf{X}}^{i}_{l}=(\hat{\bm{x}}^{i}_{l-1},\hat{\bm{x}}^{i}_{l},\hat{\bm{x}}^{i}_{l+1})$, which is then by the denoiser selected for channel $l$ . Specific to edge channels, the processes for the first and last channels form image by communicating with their only neighbor, resulting in $\hat{\mathbf{X}}^{i}_{1}=(\hat{\bm{x}}^{i}_{1},\hat{\bm{x}}^{i}_{1},\hat{\bm{x}}^{i}_{2})$ and $\hat{\mathbf{X}}^{i}_{L}=(\hat{\bm{x}}^{i}_{L-1},\hat{\bm{x}}^{i}_{L},\hat{\bm{x}}^{i}_{L})$, respectively. Thereafter, the denoisers for each channel are applied simultaneously to . To exploit the transformation invariance of images, random $90^{\circ}$ rotations and flips, and corresponding inverse transformations, are applied to the inputs and outputs of the DNNs, respectively (M. Terris et al., 2023). The result of the step all channels forms the image cube update $\mathbf{X}^{i+1}$.
IV.3.3 Dynamic denoiser selection and spectral index map update
is then monitored to decide whether the current denoiser $\operatorname{D}_{\sigma}$ and its corresponding scaling factor $\varrho_{l}$ should be updated.
We adopt a similar strategy to update the estimate of the spectral index map $\widetilde{\bm{\alpha}}$ from the current estimate of the cube $\mathbf{X}^{i}$, when required. Ignoring the spectral curvature term in (1), setting $\nu_{0}=1$ for simplicity and applying logarithm to both sides, the pixel $n$ in channel $l$ should follow
@@ IV.3.3 Dynamic denoiser selection and spectral index map update
The slope of this linear can be evaluated pixel-wise using linear regression estimate of the spectral index map $\widetilde{\bm{\alpha}}$. In practice, a smoothing 2D Gaussian kernel is applied to the image cube before linear regression numerical stability. The 2D Gaussian kernel is fitted to the PSF’s primary lobe at the highest-frequency channel, which represents the highest spatial resolution of the measurements. Additionally, pixels with intensities below a preset threshold, taken as the minimal heuristic noise level of the entire image cube, are excluded from the estimation of $\widetilde{\bm{\alpha}}$.
IV.3.4 Fully parallelized implementation
HyperAIRI reconstruction is executed in a distributed, channel-wise fashion, together with an efficient large-scale RI measurement operator implementation (A. Dabbech et al., 2025). During initialization, each process is assigned its channel’s back-projected dirty image, local measurement operator, selected denoiser and initial spectral-index map. At each FB iteration, gradient-descent and denoising steps then run concurrently across channels. Each channel process exchanges only its immediate neighbors’ images to assemble the three-slice input cubes between these two steps, minimizing communication overhead. As with the AIRI denoiser, the HyperAIRI denoiser has spatial faceting functionality. The denoisers can be applied on the overlapped facets extracted from $\hat{\mathbf{X}}^{i}_{l}$ in parallel. This hybrid spatial-spectral parallelization scales nearly linearly with both the number of channels and available compute resources. As a result, our implementation delivers a highly efficient solver capable of handling large-scale hyperspectral datasets in practical runtimes.
V Validation on simulated VLA observations